A care price is $12,000 & its market value after 7-years is $5,000. In addition its
price is ascending by 4% yearly, estimate the following:
a. Depreciation rate in the first five years of its life.
b. Total depreciation in that period also.
c. The inflation rate.
1
Expert's answer
2020-05-26T11:03:45-0400

a) "S^*=(((((((S*i)-P)*i-P)*i-P)*i-P)*i-P)*i-P)*i-P"

"S^*=S*i^7-P*\frac{i^7-1}{i-1}"

"P=(S*i^7-S^*)*\frac{i-1}{i^7-1}"

P-annual payments

S-the care price

S^*-the market value after 7-years

"i=1+0.04=1.04"

"P=12,000*1.04^7-5,000)*\frac{1.04-1}{1.04^7-1}=1,366.27"

"D=\frac{P}{S}*100\%"

D-Depreciation rate

"D=\frac{1,366.27}{12,000}*100\%=11.39\%"


b) "1,366.27*7=9563.89"

c) The inflation rate is 4%